# Saliency-Guided DETR for Moment Retrieval and Highlight Detection

### v1

@@ Abstract
Existing approaches for video moment retrieval and highlight detection are not able to align text and video features efficiently, resulting in unsatisfying performance and limited production usage. To address this, we propose a novel architecture that utilizes recent foundational video models designed for such alignment. Combined with the introduced Saliency-Guided Cross Attention mechanism and a hybrid DETR architecture, our approach significantly enhances performance in both moment retrieval and highlight detection tasks. For even better improvement, we developed InterVid-MR, a large-scale and high-quality dataset for pretraining. Using it, our architecture achieves state-of-the-art results on the QVHighlights, Charades-STA and TACoS benchmarks. The proposed approach provides an efficient and scalable solution for both zero-shot and fine-tuning scenarios in video-language tasks. The dataset and code will be published as open-source [^1] https://github.com/ai-forever/sg-detr.
@@ 1 Introduction
The task of searching for specific moments in a video based on a text query has always been in demand. With the rapid growth of video content on all platforms, it has become increasingly important. At the same time, despite their improved quality, existing solutions still have very limited performance and, therefore, limited real-world application.
@@
Existing methods mostly utilize frozen encoders, such as SlowFast [7] and CLIP [28], to extract features from videos and text queries in the QVHighlights[16] benchmark, but they often fail to align these features effectively. Some works, such as [22], have attempted to address CLIP’s limitations and enhance the quality of video features and have even achieved some performance improvement. However, recent advances in foundational video models have enabled the use of more powerful encoders for both text and video, paving the way for novel architectural approaches in the moment retrieval field. We propose a new model block that utilizes the advantages of the InternVideo2 [42] aligned video-text encoder. Our approach starts with generating initial highlight scores termed as Local Saliency Scores by directly comparing text and video embeddings and then refining them by incorporating the global context of the entire video.
@@
Inspired by the success of DETR in 2D object detection [5], the authors of the QVHighlights benchmark adapted its principles for 1D video moment localization task. The resulting model, Moment-DETR [16], became a foundational development in the field, paving the way for further works [25, 24, 15, 36]. However, other architecture variants have also been proposed to address this problem. For instance, the authors of [19] avoid the one-to-one matching typical for DETR-based models, and instead tackle moment localization using standard detection heads with one-to-many matching. In our work, we combine both strategies. Inspired by advancements in 2D detection [49], we propose a hybrid architecture for detecting relevant intervals, which incorporates both a convolutional detector head and a transformer decoder block with a one-to-one matching scheme. Additionally, we use an IoU scoring mechanism to improve the localization of video spans further.
As was claimed in [24], not all video content may be relevant to the textual query. To mitigate this issue and incorporate relevant part of textual information into the video, we propose a Saliency-Guided Cross Attention mechanism for cross-modal interaction, where attention scores are weighted with Local Saliency Scores.
@@
Another area of research in the field is the study of the interrelationship between the tasks of moment retrieval and highlight detection. Recently, the authors of TR-DETR [36] highlighted the need to ensure interaction between the heads solving these two tasks. We build upon these ideas and further integrate the tasks of moment retrieval and highlight detection using our Saliency Amplifier module.
@@
Despite the high quality of the QVHighlights dataset’s annotations, its limited size necessitates additional data for effective model pretraining. To address this, we introduce a novel data generation approach that utilizes large video retrieval datasets, such as InterVid-FLT-10m [41], and leverages the impressive capabilities of foundational text-video models to create high-quality training data. This method enables us to build a pretraining dataset called InterVid-MR, which contains 150k samples. Our model, trained solely on the InterVid-MR dataset, demonstrates strong performance in zero-shot evaluations on the QVHighlights benchmark. Moreover, fine-tuned on QVHighlights data, the model achieves state-of-the-art results.
The contributions of our work can be summarized as:
@@
  • We conducted extensive experiments to evaluate the effectiveness of the proposed model across three benchmarks: QVHighlights, Charades-STA and TACoS.
@@ 2 Related Works
Moment retrieval (MR) aims to localize specific video intervals based on a given textual query. Approaches for this task are typically classified into proposal-based and proposal-free methods. Proposal-based methods utilize predefined proposals, such as sliding windows [1, 10] or temporal anchors [6]. These methods often require complex pre- and post-processing, with performance being heavily influenced by the quality of the generated proposals. In contrast, proposal-free methods [16, 19, 24] leverage multimodal information and directly predict the start and end coordinates of the target moments, thus bypassing the need for explicit proposal generation.
Highlight detection (HD) focuses on assigning a relevance score to each video sub-clip based on the textual query, allowing for easy ranking. Methods in this domain are typically divided into three categories based on the level of annotation: supervised [12, 37], which requires precise and often expensive annotations; weakly supervised [4, 45], which learns to identify key moments using event labels; and unsupervised [2, 23, 32], which operates without the need for labeled data.
@@
Research in multimodal alignment, which integrates visual and textual modalities, has shown significant potential for improving performance in downstream tasks such as moment retrieval and highlight detection. Various methods have been proposed to tackle this challenge.
@@
Figure 1: Overall architecture of our framework. Detailed explanations of notations are described in Sec. 3
3 Model
Moment retrieval and highlight detection tasks aim to identify key moments in a video based on a textual query. In this section, we describe in detail each component of our proposed architecture, which effectively solves both tasks. An overview of the proposed Saliency-Guided DETR (SG-DETR) is shown in Figure 1.
@@ 3.1 Feature Extraction
Considering the issues mentioned above, we decided to use the InternVideoV2 [42] model, which was trained for aligning video and text. It effectively handles the modeling of spatial-temporal relationships and does not require the inclusion of additional models like SlowFast in the pipeline, significantly simplifying it.
Our feature extraction pipeline can be described as follows. We divide video into 2 seconds clips and extract features for each clip using the video encoder of the InternVideoV2-1B model: $F_{v}=[f_{v}^{1},f_{v}^{2},\ldots,f_{v}^{L}]\in\mathbb{R}^{L\times d_{v}}$, where $L$ and $d_{v}$ are the number of clips and the visual feature dimension, respectively. Similarly, text features are extracted using the text encoder of the InternVideoV2-1B model: $F_{t}=[f_{t}^{1},f_{t}^{2},\ldots,f_{t}^{N}]\in\mathbb{R}^{N\times d_{t}}$, where $N$ and $d_{t}$ are the number of the textual tokens and the textual feature dimension, respectively.
@@ 3.2 Local Saliency Scores
where PoolingEncoder is an encoder consisting of $N$ layers of AttentionPooling modules, and $F_{\text{sent}}\in\mathbb{R}^{1\times d_{t}}$ represents the sentence token.
@@ 3.3 Saliency-Guided Cross Attention
As it was highlighted in the CG-DETR paper [24], not all video content aligns with the semantics of a text query. However, when softmax is used as the key element of cross-attention, it can result in assigning similar importance to all query tokens even for irrelevant ones. This behavior can reduce the accuracy of the model in identifying the true correspondence between the video content and the text query. Although switching from softmax to sigmoid seems like a straightforward solution, it introduces limitations by removing the dependencies between tokens, which weakens the ability to rank text-relevance scores accurately.
To address these issues, we introduce the Saliency-Guided Cross Attention (SGCA) mechanism, which effectively overcomes the limitations of standard cross-attention by considering the local relevance of each individual video clip to the text query. Formally, let the query tokens be derived from the video tokens as $Q=[p_{Q}(f_{v}^{1}),p_{Q}(f_{v}^{2}),\ldots,p_{Q}(f_{v}^{L})]$, key and value tokens be derived from the text tokens: $K=[p_{K}(f_{t}^{1}),p_{K}(f_{t}^{2}),\ldots,p_{K}(f_{t}^{L})]$ and $V=[p_{V}(f_{t}^{1}),p_{V}(f_{t}^{2}),\ldots,p_{V}(f_{t}^{L})]$, where $p_{Q}(\cdot)$, $p_{K}(\cdot)$, and $p_{V}(\cdot)$ are learned projection functions that map the video and text tokens to a common hidden dimension $h$.
@@
The described attention mechanism is then used as a part of the Cross-Modal Transformer Encoder to derive the video token representation enriched with textual information.
@@
After that, we apply a standard Transformer Encoder consisting of N layers to incorporate global context into features.
$\widehat{{F_{v}}^{T}}=\text{TransformerEncoder}({F_{v}}^{T})$ (6)
@@ 3.4 From Local to Global Saliency scores
$S_{\text{global}}=S_{\text{local}}+\text{Linear}(\widehat{{{F_{v}}^{T}}_{l}}).$ (8)
We reapply the SA module once the global saliency scores are obtained. This time, the goal is to leverage the information obtained during the Highlight Detection task to address the Moment Retrieval task, as we now know which features are more relevant and which are less.
@@
We sent globally amplified video features $\widehat{{{F_{v}}^{T}}_{g}}$ to regression heads.
@@ 3.5 Hybrid Detector
In previous works, various detection heads, including CNN-based [19] and DETR-like models [25, 24, 16], have tackled the moment retrieval task. Co-DETR [49] has demonstrated that combining these approaches can be adequate for object detection tasks in the image domain. Inspired by this, we employ a hybrid detector with two detection heads. The ATSS [47] detector serves as the auxiliary CNN-based head, while the DINO-DETR [46] detector is used as the primary detection head. In line with Co-DETR, we utilize positive anchors from the ATSS detector as queries for the DETR head during model training. Consequently, during training, the DETR head receives three sets of references: a few groups of noisy target spans, positive anchors from ATSS, and primary references generated by the query selector mechanism introduced in DINO-DETR. As part of the post-processing stage, we follow [34] and apply a weighted fusion of ATSS and DETR boxes to produce the final predictions.
@@ 3.6 Localization confidence
Prior research, including works such as [14, 44, 9], has demonstrated that predicting IoU scores or a similar localization confidence metric significantly improves object detection performance. Building on these insights, we introduce an additional DETR head dedicated to predicting IoU scores between the predicted and ground truth spans.
@@
$S_{det}=p_{i}^{\alpha}\hat{IoU_{i}}^{(1-\alpha)}$ (10)
where $p_{i}$ is the classification confidence and $\hat{IoU_{i}}$ is the predicted IoU.
3.7 Training Objectives
Our model is trained using losses categorized into three main groups.
Highlight Detection Task
Moment Retrieval Task
For the moment retrieval task, we use CE loss, generalized IoU (GIoU) loss [31], and smooth L1 loss to train both DETR and ATSS detection heads. For the ATSS head [47], we also incorporate Centerness Loss [39] to enhance localization precision. In the DETR head, CE loss is applied to IoU head predictions. Losses for auxiliary DETR queries are computed similarly to primary queries. The objectives for the task are:
Auxiliary Losses
$\mathcal{L}_{aux}=\mathcal{L}_{bce}+\mathcal{L}_{align}$ (14)
Overall Objective
@@ 4 Experiments and Results
TR-DETR InternVideo2-1b 70.08 $\pm$ 0.15 49.20 $\pm$ 0.50 47.99 $\pm$ 0.42 43.43 $\pm$ 0.16 71.13 $\pm$ 0.25
Ours InternVideo2-1b 73.52 $\pm$ 0.05 57.91 $\pm$ 0.13 55.64 $\pm$ 0.20 43.91 $\pm$ 0.14 71.47 $\pm$ 0.73
@@ 4.1 Datasets
QVHighlights is the primary benchmark annotated for both moment retrieval and highlight detection tasks. The dataset is relatively small and consists of approximately 10k videos with human-written text queries covering various topics, from everyday activities and travel in lifestyle vlog videos to social and political activities in news videos. Charades-STA is a dataset used exclusively for moment retrieval task evaluation. It includes 16k sentence-moment pairs from 10k videos of indoor activities with an average duration of 30 seconds. TACoS consists of 127 cooking videos with an average duration of 287 seconds and 19k sentence-moment pairs. This dataset is recognized as particularly challenging because the moments occupy only a tiny portion of the considerably long videos.
@@ 4.2 Feature Extraction
In earlier studies [21, 16, 25, 13], various feature extractors were used on the mentioned benchmarks. For the QVHighlights benchmark, a combination of CLIP [28] and SlowFast [7] models was employed. The Charades-STA benchmark used the VGG [33] extractor for video content and GLOVE [27] embeddings for text queries. This variety in methods requires maintaining numerous feature extraction pipelines and tuning hyperparameters for each, which slows down research and increases the carbon footprint. Additionally, many of these methods have become outdated and hardly remain representative. We propose using a unified approach to overcome the problem. Specifically, we extracted visual and textual features for each benchmark using only the InterVidV2-1b [42] model. We retrained the primary existing models using the InterVidV2-1b embeddings to assess the new features’ impact and validate our method’s competitiveness on QVHighlights validation set. The results are presented in Tab. 1.

### v2

@@ Abstract
With the rapid growth of video content available, the ability to search for specific moments within videos using textual queries has become increasingly relevant. This is crucial in many scenarios, from surveillance cameras where it may be necessary to find specific events in extensive video streams to searching for exciting movie scenes. However, existing approaches for video Moment Retrieval and Highlight Detection often struggle to effectively align text and video features, limiting their performance. We argue that utilizing recent foundational video models designed for video-text alignment can overcome these limitations. We propose a novel architecture that utilizes such models to test this hypothesis. Combined with our novel Saliency-Guided Cross Attention mechanism and a hybrid DETR architecture, our approach provides significantly improved results. To further enhance our approach, we developed InterVid-MR — a large-scale, high-quality dataset specifically designed for a pretraining stage. Extensive experiments and comparisons with current state-of-the-art methods confirm the effectiveness of the approach, achieving 58.8 mAP on QVHighlights, 60.7 R@1 mIoU on Charades-STA, and 42.4 R@1 mIoU on TACoS. These results highlight the efficiency and scalability of the method for video-language tasks in both zero-shot and fine-tuning scenarios. The dataset and code will be publicly available.  [^1] https://gracikk-ds.github.io/sg-detr/
@@ 1 Introduction
The task of searching for specific moments in a video based on a text query has always been in demand. With the rapid growth of video content, it has become increasingly relevant. For instance, in the entertainment industry, quickly locating scenes of interest like epic battles in movies or specific landmarks in travel videos can greatly enhance content accessibility and user engagement. In sports analytics, identifying key events can be very important for streaming platforms and spectators. Similarly, in security surveillance, rapid search of incidents like car accidents, crime cases, or safety violations in industrial settings is vital for timely response or investigations.
@@
Despite this demand, existing solutions for video moment retrieval (MR), which aims to localize specific video intervals, and highlight detection (HD), which focuses on assigning a relevance score to each video sub-clip, exhibit limited performance and scalability, constraining their practical adoption. Most of the methods utilize frozen encoders, such as SlowFast [8] and CLIP [29], to extract features from videos and text queries in the QVHighlights [17] benchmark, but they often fail to align these features effectively. Some works, such as [21], have attempted to address CLIP’s limitations and enhance the quality of video features and have even achieved some performance improvement. However, recent advances in foundational video models have enabled the use of more powerful encoders for both text and video, paving the way for novel architectural approaches in the moment retrieval field. We propose a new model block that utilizes the advantages of the InternVideo2 [42] aligned video-text encoder. Our approach starts with generating initial highlight scores termed as Local Saliency Scores by directly comparing text and video embeddings and then refining them by incorporating the global context of the entire video.
@@
Inspired by the success of DETR in 2D object detection [6], the authors of the QVHighlights benchmark adapted its principles for 1D video moment localization task. The resulting model, Moment-DETR [17], became a foundational development in the field, paving the way for further works [26, 25, 16, 37]. However, other architecture variants have also been proposed to address this problem. For instance, the authors of [19] avoid the one-to-one matching typical for DETR-based models, and instead tackle moment localization using standard detection heads with one-to-many matching. Inspired by advancements in 2D detection [50], we combine both strategies and propose a hybrid architecture for detecting relevant intervals, which incorporates a convolutional detector head and a transformer decoder block with a one-to-one matching scheme. Additionally, we use an IoU scoring mechanism to improve the localization of video spans further.
Different parts of a video vary in their relevance to a text query. To focus the model on the most relevant content, we propose a Saliency-Guided Cross Attention (SGCA) mechanism for cross-modal interaction, where cross-attention scores are weighted by the Local Saliency Scores (Sec. 3.2). By avoiding extra tokens or complex gating modules, SGCA provides a simpler implementation compared to prior approaches [25, 37].
@@
Another area of research focuses on the interrelationship between the tasks of moment retrieval and highlight detection. Recently, the authors of TR-DETR [37] emphasized the need to ensure interaction between the heads solving these two tasks. We build upon these ideas and further integrate the tasks of moment retrieval and highlight detection using Saliency Amplifier module.
@@
Despite the high quality of the QVHighlights dataset’s annotations, its limited size necessitates additional data for effective model pretraining. To address this, we introduce a novel data generation approach that utilizes large video retrieval datasets, such as InterVid-FLT-10m [43], and leverages the impressive capabilities of foundational text-video models to create high-quality training data. This method enables us to build a pretraining dataset called InterVid-MR, which contains 150k samples. The model, trained solely on the InterVid-MR dataset, demonstrates strong performance in zero-shot evaluations on the QVHighlights benchmark. Moreover, fine-tuned on QVHighlights data, the model achieves state-of-the-art results.
The contributions of the work can be summarized as:
@@
  • We conducted extensive experiments and achieved state-of-the-art results across three benchmarks: QVHighlights, Charades-STA, and TACoS.
@@ 2 Related Works
Approaches for moment retrieval task are typically classified into proposal-based and proposal-free methods. Proposal-based methods utilize predefined proposals, such as sliding windows [1, 11] or temporal anchors [7]. These methods often require complex pre- and post-processing, with performance being heavily influenced by the quality of the generated proposals. In contrast, proposal-free methods [17, 19, 25] leverage multimodal information and directly predict the start and end coordinates of the target moments, thus bypassing the need for explicit proposal generation.
Methods for the highlight detection task are typically divided into three categories based on the level of annotation: supervised [13, 38], which requires precise and often costly annotations; weakly supervised [4, 46], which learns to identify key moments using event labels; and unsupervised [2, 24, 33], which operates without labeled data.
@@
Research in multimodal alignment, which integrates visual and textual modalities, has shown significant potential for improving performance in downstream tasks such as moment retrieval and highlight detection. Various methods have been proposed to tackle this challenge [43, 42].
@@
Figure 1: Saliency-Guided DETR architecture. Detailed explanations of notations are described in Sec. 3.
3 Method
In this section, we describe each component of the proposed architecture, which effectively solves both tasks. An overview of the proposed Saliency-Guided DETR (SG-DETR) is shown in Fig. 1.
@@ 3.1 Feature Extraction
Considering the abovementioned issues, we decided to use the InternVideoV2 [42] model, which was trained for aligning video and text. It effectively handles the modeling of spatial-temporal relationships and does not require the inclusion of additional models like SlowFast in the pipeline, significantly simplifying it.
Our feature extraction pipeline can be described as follows. We divide the video into 2-second clips. Then, we extract text and video features using the InternVideoV2-1B model. These features are then projected to a new dimension $d$: $F_{v}=[f_{v}^{1},f_{v}^{2},\ldots,f_{v}^{L}]\in\mathbb{R}^{L\times d}$ for video and $F_{t}=[f_{t}^{1},f_{t}^{2},\ldots,f_{t}^{N}]\in\mathbb{R}^{N\times d}$ for text, where $L$ and $N$ denote the number of clips and textual tokens, respectively.
@@ 3.2 Local Saliency Scores
where PoolingEncoder is an encoder consisting of $N$ layers of AttentionPooling modules, and $F_{\text{sent}}\in\mathbb{R}^{1\times d}$ represents the sentence token.
@@ 3.3 Saliency-Guided Cross Attention
In standard cross-attention, softmax forces each video token to absorb some textual information, even when the correspondence is weak or irrelevant. As a result, video segments that don’t match the text query may still be influenced by it, degrading the model’s ability to distinguish relevant and irrelevant segments precisely. A possible solution is to replace softmax with sigmoid activations. However, using sigmoid breaks a text query’s inter-token dependency, weakening the model’s ability to rank relevant content [25].
To address this issue, we introduce the Saliency-Guided Cross Attention (SGCA) mechanism, which effectively overcomes the limitations of standard cross-attention by considering the local relevance of each individual video clip to the text query. Formally, let the query tokens be derived from the video tokens as $Q=[p_{Q}(f_{v}^{1}),p_{Q}(f_{v}^{2}),\ldots,p_{Q}(f_{v}^{L})]$, key and value tokens be derived from the text tokens: $K=[p_{K}(f_{t}^{1}),p_{K}(f_{t}^{2}),\ldots,p_{K}(f_{t}^{L})]$ and $V=[p_{V}(f_{t}^{1}),p_{V}(f_{t}^{2}),\ldots,p_{V}(f_{t}^{L})]$, where $p_{Q}(\cdot)$, $p_{K}(\cdot)$, and $p_{V}(\cdot)$ are learned projection functions that map the video and text tokens to a common hidden dimension $h$.
@@
The proposed SGCA mechanism imposes only a minor computational burden, as detailed in Appendix A.
The described attention mechanism is then used as a part of the Cross-Attention Transformer to derive the video token representation enriched with textual information. After that, the standard Transformer Encoder is applied [26].
@@
$\widehat{{F_{v}}^{T}}$ $=\text{TransformerEncoder}({F_{v}}^{T})$
@@ 3.4 From Local to Global Saliency scores
[…]
We reapply the SA module once the global saliency scores are obtained. The goal is to leverage the information obtained during the HD task to address the MR task.
@@
We sent globally amplified features $\widehat{{{F_{v}}^{T}}_{g}}$ to regression heads.
@@ 3.5 Hybrid Detector
In previous works, various detection heads, including CNN-based [19] and DETR-like models [26, 25, 17], have tackled the MR task. Co-DETR [50] has demonstrated that combining these approaches can be adequate for object detection tasks in the image domain. Inspired by this, we employ a hybrid detector with two detection heads. The ATSS [48] detector serves as the auxiliary CNN-based head, while the DINO-DETR [47] detector is used as the primary detection head. In line with Co-DETR, we utilize positive anchors from the ATSS detector as queries for the DETR head during model training. Consequently, during training, the DETR head receives three sets of references: a few groups of noisy target spans, positive anchors from ATSS, and primary references generated by the query selector mechanism introduced in DINO-DETR. As part of the post-processing stage, we follow [35] and apply a weighted fusion of ATSS and DETR boxes to produce the final predictions.
@@ 3.6 Localization confidence
Prior research, including works such as [15, 45, 10, 5], has demonstrated that predicting IoU scores or a similar localization confidence metric significantly improves object detection performance. Building on these insights, we introduce an additional DETR head dedicated to predicting IoU scores between the predicted and ground truth spans.
@@
$S_{det}=p_{i}^{\gamma}\hat{IoU_{i}}^{(1-\gamma)}$ (9)
where $p_{i}$ is the classification confidence, $\hat{IoU_{i}}$ is the predicted IoU, and $\gamma$} is a hyperparameter that controls the balance between $p_{i}$ and $\hat{IoU_{i}}$.
@@ 4 Experiments and Results
InternVideo2-1b 70.08 $\pm$ 0.15 49.20 $\pm$ 0.50 47.99 $\pm$ 0.42 43.43 $\pm$ 0.16 71.13 $\pm$ 0.25
SG-DETR InternVideo2-1b 73.52 $\pm$ 0.05 57.91 $\pm$ 0.13 55.64 $\pm$ 0.20 43.91 $\pm$ 0.14 71.47 $\pm$ 0.73
@@ 4.1 Datasets
QVHighlights is the primary benchmark annotated for both MR and HD tasks. The dataset is relatively small and consists of approximately 10k videos with human-written text queries covering various topics, from everyday activities and travel in lifestyle vlog videos to social and political activities in news videos. Charades-STA is a dataset used exclusively for moment retrieval task evaluation. It includes 16k sentence-moment pairs from 10k videos of indoor activities with an average duration of 30 seconds. TACoS consists of 127 cooking videos with an average duration of 287 seconds and 19k sentence-moment pairs. This dataset is recognized as particularly challenging because the moments occupy only a tiny portion of the considerably long videos.
@@ 4.2 Feature Extraction
In earlier studies [22, 17, 26, 14], various feature extractors were used on the mentioned benchmarks. For the QVHighlights benchmark, a combination of CLIP [29] and SlowFast [8] models was employed. The Charades-STA benchmark used the VGG [34] extractor for video content and GLOVE [28] embeddings for text queries. This variety in methods requires maintaining multiple feature extraction pipelines and tuning hyperparameters for each, which slows research and increases the carbon footprint. Many of these methods have also become outdated and hardly remain representative. We propose a unified approach to address the problem. Specifically, we extracted visual and textual features for each benchmark using only the InterVidV2-1b [42] model. We retrained some existing models using the InterVidV2-1b embeddings to assess the new features’ impact and validate our method’s competitiveness on the QVHighlights validation set. Although CLIP+SlowFast features allow additional comparisons with prior methods, they lack the necessary alignment for our SGCA module’s text–video similarity computation, leading to unstable training and reduced performance. The results are presented in Tab. 1.